# Three-dimensional Dynamics of Strongly Magnetized Ion–Electron Relativistic Reconnection

### v1

@@ 1 Introduction
Reconnection in semirelativistic ion–electron plasmas has so far received less attention than both the nonrelativistic and (ultra)relativistic regimes, although understanding the relative energization of electrons and ions is of critical importance for global magnetohydrodynamic (MHD) modeling of radiation from black-hole accretion flows (e.g. Chael et al., 2018; Dexter et al., 2020; Ressler et al., 2020; Hankla et al., 2022; Scepi et al., 2022; Ressler et al., 2023). Unlike in the (ultra)relativistic regime where ions behave essentially as positrons yielding symmetric behavior between electrons and ions (unless radiative effects are considered; e.g. Chernoglazov et al., 2023), in the semirelativistic regime electrons and ions at the same temperature have different gyroradii and behave differently as they are energized. Important progress has been made through 2D PIC studies of reconnection in this regime (Melzani et al., 2014a, b; Guo et al., 2016; Rowan et al., 2017; Ball et al., 2018b; Werner et al., 2018; Ball et al., 2019; Rowan et al., 2019). However, there remains much we do not understand about reconnection and other dissipation mechanisms in 3D current sheets. In relativistic electron–positron plasma, some earlier PIC studies concluded that the 3D behavior of initially stable current sheets is fairly similar to the 2D dynamics (Yin et al., 2008; Liu et al., 2011; Kagan et al., 2013; Guo et al., 2014; Sironi & Spitkovsky, 2014; Guo et al., 2015; Werner & Uzdensky, 2017; Guo et al., 2021), although there is significant evidence that, at the very least, particle energy in 2D is limited by the difficulty of escape from plasmoids formed by reconnection, whereas in 3D particles escape and undergo additional acceleration (e.g. Dahlin et al., 2015; Petropoulou & Sironi, 2018; Li et al., 2019; Hakobyan et al., 2021; Zhang et al., 2021a, 2023). In 2D collisionless reconnection, dynamics is restricted to the plane where tearing modes can develop and grow. Plasmoids can form from reconnection, and in periodic or closed domains, they accumulate and create system-size-scale structures where particles are trapped indefinitely. This slows down (or halts) reconnection, preventing further dissipation of magnetic energy. In 3D domains, additional out-of-plane modes such as the drift-kink instability (DKI) and the flux-rope (MHD) kink instability may develop, folding and rippling current sheets perpendicularly to the reconnection plane. The DKI in particular is stronger for higher magnetizations and for plasmas with small scale separation between species, (e.g. pair plasma, or the semirelativistic ion–electron plasma considered in this work), and may nonlinearly interact with tearing modes. MHD kinking, instead, can destroy elongated magnetic structures (plasmoids in 2D) created by reconnection, liberating trapped particles and contributing to their energization. In some cases, these kinking modes can alter the mechanism and nature of magnetic energy dissipation and particle energization (e.g. Pritchett et al., 1996; Zenitani & Hoshino, 2005a, 2007, 2008; Hoshino, 2020; Werner & Uzdensky, 2021; Hoshino, 2024). These differences have been found to carry over to the 3D semirelativistic regime (Werner & Uzdensky, 2024), but very limited exploration of 3D dynamics has been carried out so far. In particular, it is unclear how previous results (focusing on low magnetization) extend to strongly magnetized regimes.
@@ 3.1 Magnetic-energy Dissipation
In our runs, we can generally identify an initial “quiet” phase during which the initial equilibrium remains marginally stable and (noise-induced) perturbations start growing. Over time, these fluctuations reach amplitudes large enough to induce reconnection and fragment the initial current sheet. At that point, magnetic energy is violently released through an “active” reconnection phase; a subsequent, final phase is then initiated, whose dynamics depends on upstream conditions, as we discuss below. In general, during active reconnection, we can identify the presence of flux tubes extending along $z$ and persistent chaoticity in the current-sheet structure. Toward the end of the simulation, the current sheet becomes thick and rather unstructured, without visible macroscopic flux tubes. This dynamics qualitatively agrees with the results of Werner & Uzdensky (2021, 2024).
@@
3.2 3D Dynamics: Tearing versus Kink Modes
@@
[…]
At low $\sigma_{i,0}$ (panels $a$ and $d$), tearing modes start to develop much faster than kink modes, implying that the current sheet experiences thinning and breaking in a 2D-like fashion, before significant kink fluctuations develop. Power in tearing modes peaks, and shortly after that kink modes reach saturation. Tearing modes then start decreasing in power, while kink modes remain active. It is noteworthy that the only significant kink modes are those at the smallest $k$ (i.e. at wavelengths comparable to the system size), implying that MHD-kink is active but DKI is not. This aligns with linear theory, which predicts that DKI is rather insignificant at low magnetization (e.g. Zenitani & Hoshino 2005b).
@@
Figure 3: Volume rendering of the total (ion and electron) plasma density in the current sheet during subsequent reconnection phases (from top to bottom) and for increasing magnetization (from left to right).
@@ 3.3 Current-sheet Structure and Kinking Flux Tubes
These results indicate that DKI-driven sheet thickening, which occurs faster and more strongly at high $\sigma_{i,0}$, is the key mechanism driving qualitatively different reconnection dynamics in our 3D simulations. This is further supported by Fig. 4, where we show the evolution of the average current-sheet thickness $\delta$, defined as the region where $\sigma_{i}\leq 3\sigma_{i,0}/4$, i.e. where the magnetization drops by at least 25% from its upstream value. We observe that the sheet becomes much thicker at a much faster rate for higher magnetizations. This again suggests that DKI dynamics changes the sheet structure at high $\sigma_{i,0}$, maintaining the sheet in a chaotic, increasingly thicker state in which tearing modes (and the creation of large-scale flux ropes) are strongly suppressed.
@@ 3.4 Particle acceleration
Our ion–electron simulations allow us to investigate how different plasma species can gain energy via reconnection. To conclude our analysis, Fig. 5 shows the energy distributions $f(\gamma)\equiv\mathrm{d}\mathrm{N}/\mathrm{d}\gamma$ of electrons (top row) and ions (bottom row) for equivalent 2D (left column) and 3D (right column) simulations, for $\sigma_{i,0}\in[0.1,10]$ and at the end of each run ($t_{\mathrm{f}}=2000\omega_{\mathrm{p},i}^{-1}$). The initial distribution at $t=0$ is shown as a dashed black line. The trend we retrieve corresponds to the general expectation that higher $\sigma_{i,0}$ produce higher-energy particles (e.g. Werner & Uzdensky 2021), but with important differences between species and between 2D and 3D runs.
@@
First, we note that both particle species can develop nonthermal populations, with $\sigma_{i,0}=10$ producing evident nonthermal tails extending for up to 2 decades in both 2D and 3D. However, the slope of the nonthermal tail differs depending on dimensionality, with 3D runs producing markedly less steep tails. Reference slopes of $\gamma^{-1.5}$, $\gamma^{-2}$, and $\gamma^{-2.5}$ are shown to guide the eye; this broadly aligns with previous findings for 3D reconnection (Zhang et al. 2021a, 2023). In these runs, ions are initially nonrelativistic and, for the highest $\sigma_{i,0}$, can reach maximum Lorentz factors of 10similar-toabsent10$\sim$ 10∼ 10100100100100; electrons are already relativistic at initialization, and their energy gain is such that $\gamma$ can increase by over 2 orders of magnitude (up to $\sim 10^{5}$) for the case with the highest magnetization. The maximum Lorentz factor is roughly comparable in 2D and 3D for both species, but in 3D it is generally smaller by a factor 2–3 for the highest $\sigma_{i,0}$.
@@ 4 Discussion
[…]
These findings contradict the standard picture in which 3D reconnection dissipates more magnetic energy than in 2D (e.g. Werner & Uzdensky 2021; Zhang et al. 2021b; Werner & Uzdensky 2024). In particular, our results describe a scenario in which reconnecting plasma gathers in slowly dissipating current sheets, which are kept thick and chaotic by DKI. One such current sheet may even be described as two separate sheets, where upstream, cold- and magnetized-plasma conditions meet an internal hot, lowly magnetized plasma state. In this situation, standard expectations for symmetric Harris-sheet reconnection may not be applicable. Our high-$\sigma_{i,0}$ results are especially relevant for astrophysical scenarios such as compact-object magnetospheres and coronae, where ultramagnetized plasma could be undergoing reconnection, potentially powering high-energy flares. This may have implications for comparing the time scales of reconnection with flaring activity observed from high-energy sources, although quantifying the impact of the current-sheet structure on flare dynamics requires dedicated efforts to be pursued elsewhere.

### v2

@@ 1 Introduction
Reconnection in semirelativistic ion–electron plasmas has so far received less attention than both the nonrelativistic and (ultra)relativistic regimes, although understanding the relative energization of electrons and ions is of critical importance for global magnetohydrodynamic (MHD) modeling of radiation from black-hole accretion flows (e.g. Chael et al., 2018; Dexter et al., 2020; Ressler et al., 2020; Hankla et al., 2022; Scepi et al., 2022; Ressler et al., 2023; Vos et al., 2024). Unlike in the (ultra)relativistic regime where ions behave essentially as positrons yielding symmetric behavior between electrons and ions (unless radiative effects are considered; e.g. Chernoglazov et al., 2023), in the semirelativistic regime electrons and ions at the same temperature have different gyroradii and behave differently as they are energized. Important progress has been made through 2D PIC studies of reconnection in this regime (Melzani et al., 2014a, b; Guo et al., 2016; Rowan et al., 2017; Ball et al., 2018b; Werner et al., 2018; Ball et al., 2019; Rowan et al., 2019). However, there remains much we do not understand about reconnection and other dissipation mechanisms in 3D current sheets. In relativistic electron–positron plasma, some earlier PIC studies concluded that the 3D behavior of initially stable current sheets is fairly similar to the 2D dynamics (Yin et al., 2008; Liu et al., 2011; Kagan et al., 2013; Guo et al., 2014; Sironi & Spitkovsky, 2014; Guo et al., 2015; Werner & Uzdensky, 2017; Guo et al., 2021), although there is significant evidence that, at the very least, particle energy in 2D is limited by the difficulty of escape from plasmoids formed by reconnection, whereas in 3D particles escape and undergo additional acceleration (e.g. Dahlin et al., 2015; Petropoulou & Sironi, 2018; Li et al., 2019; Hakobyan et al., 2021; Zhang et al., 2021a, 2023). In 2D collisionless reconnection, dynamics is restricted to the plane where tearing modes can develop and grow. Plasmoids can form from reconnection, and in periodic or closed domains, they accumulate and create system-size-scale structures where particles are trapped indefinitely. This slows down (or halts) reconnection, preventing further dissipation of magnetic energy. In 3D domains, additional out-of-plane modes such as the drift-kink instability (DKI) and the flux-rope (MHD) kink instability may develop, folding and rippling current sheets perpendicularly to the reconnection plane. The DKI in particular is stronger for higher magnetizations and for plasmas with small scale separation between species, (e.g. pair plasma, or the semirelativistic ion–electron plasma considered in this work), and may nonlinearly interact with tearing modes. MHD kinking, instead, can destroy elongated magnetic structures (plasmoids in 2D) created by reconnection, liberating trapped particles and contributing to their energization. In some cases, these kinking modes can alter the mechanism and nature of magnetic energy dissipation and particle energization (e.g. Pritchett et al., 1996; Zenitani & Hoshino, 2005a, 2007, 2008; Hoshino, 2020; Werner & Uzdensky, 2021; Hoshino, 2024). These differences have been found to carry over to the 3D semirelativistic regime (Werner & Uzdensky, 2024), but very limited exploration of 3D dynamics has been carried out so far. In particular, it is unclear how previous results (focusing on low magnetization) extend to strongly magnetized regimes.
@@ 3.1 Magnetic-energy Dissipation
In our runs, we can generally identify an initial “quiet” phase during which the initial equilibrium remains marginally stable and (noise-induced) perturbations start growing. Over time, these fluctuations reach amplitudes large enough to induce reconnection and fragment the initial current sheet. At that point, magnetic energy is violently released through an “active” reconnection phase; a subsequent, final phase is then initiated, whose dynamics depends on upstream conditions, as we discuss below. In general, during active reconnection, we can identify the presence of flux tubes extending along $z$ and persistent chaoticity in the current-sheet structure. Toward the end of the simulation, the current sheet becomes thick and rather unstructured, without visible macroscopic flux tubes. This dynamics qualitatively agrees with the results of Werner & Uzdensky (2021, 2024). As described in Section 2, all our simulations are run for the same physical time $t_{\mathrm{f}}=2000\omega_{\mathrm{p},i}^{-1}$, which for low magnetization (therefore small $L$) is equivalent to $\sim 111L/c$. A priori, it would seem unnecessary to let the system evolve over so many light-crossing times, since with the standard (collisionless) reconnection rate in the relativistic limit ($\sim 0.1c$) all magnetic energy would be depleted much faster. However, at low magnetization the reconnection rate will be far from the relativistic expectation, and therefore it becomes necessary to run for longer times.
@@
3.2 Tearing versus Kink Modes in 3D
@@
To quantitatively assess the growth of tearing and kink modes, we compute the Fourier spectrum of magnetic-field fluctuations in the current-sheet plane. This is shown for selected wavenumbers in Fig. 2 for $B_{y}(x)$ (top row, indicating tearing) and $B_{x}(z)$ (bottom row, indicating kink activity) for low ($\sigma_{i,0}=0.1$, left column), intermediate ($\sigma_{i,0}=0.4$, middle column), and high ($\sigma_{i,0}=10$, right column) magnetizations. The intermediate $\sigma_{i,0}=0.4$ case was selected as a representative of the threshold magnetization above which the energy-dissipation trend in 3D reverses (cf. Fig. 1bb). Here, we plot the first 15 wavenumbers $k=[k_{\mathrm{min}},2k_{\mathrm{min}},3k_{\mathrm{min}},$...], with darker lines corresponding to smaller wavenumbers and with $k_{\mathrm{min}}\equiv\pi/L$. Here, we focus on two key parameters: the time scales of the growth of tearing and kink modes, and the power in different modes.
At low $\sigma_{i,0}$ (panels $a$ and $d$), tearing modes begin to develop much faster than kink modes [^3] Note that panel $d$ shows the presence of kink oscillations at very early times at powers (a factor few) larger than those at which tearing modes exist. However, this power is associated with small $k$, i.e. with long wavelengths, which normally would indicate MHD-kinking. At this early stage, however, MHD-kink cannot be developing yet, and therefore we interpret this signal as some initial oscillation of the whole sheet, potentially induced by the initial conditions, without a specific physical meaning., implying that the current sheet experiences thinning and breaking in a 2D-like fashion, before significant kink fluctuations develop. Power in tearing modes peaks, and shortly after that kink modes reach saturation. Tearing modes then start decreasing in power, while kink modes remain active. It is noteworthy that the only significant kink modes are those at the smallest $k$ (i.e. at wavelengths comparable to the system size), implying that MHD-kink is active but DKI is not. This aligns with linear theory, which predicts that DKI is rather insignificant at low magnetization (e.g. Zenitani & Hoshino 2005b).
@@
Figure 3: Volume rendering (covering the entire simulation domain) of the total (ion and electron) plasma density in the current sheet during subsequent reconnection phases (from top to bottom) and for increasing magnetization (from left to right).
@@ 3.3 Current-sheet Structure and Kinking Flux Tubes
These results indicate that DKI-driven sheet thickening, which occurs faster and more strongly at high $\sigma_{i,0}$, is the key mechanism driving qualitatively different reconnection dynamics in our 3D simulations. This is further supported by Fig. 4, where we show the evolution of the average current-sheet thickness $\delta$, defined as the region where $\sigma_{i}\leq 3\sigma_{i,0}/4$, i.e. where the magnetization drops by at least 25% from its upstream value [^4] We have tested different thresholds for this analysis down to $\sigma_{i}\leq 0.5\sigma_{i,0}$, obtaining qualitatively similar results.. We observe that the sheet becomes much thicker at a much faster rate for higher magnetizations. This again suggests that DKI dynamics changes the sheet structure at high $\sigma_{i,0}$, maintaining the sheet in a chaotic, increasingly thicker state in which tearing modes (and the creation of large-scale flux ropes) are strongly suppressed.
@@ 3.4 Particle acceleration
Our ion–electron simulations allow us to investigate how different plasma species can gain energy via reconnection. To conclude our analysis, Fig. 5 shows the energy distributions $f(\gamma)\equiv\mathrm{d}\mathrm{N}/\mathrm{d}\gamma$ (excluding the initial drifting populations) of electrons (top row) and ions (bottom row) for equivalent 2D (left column) and 3D (right column) simulations, for $\sigma_{i,0}\in[0.1,10]$ and at the end of each run ($t_{\mathrm{f}}=2000\omega_{\mathrm{p},i}^{-1}$). The initial distribution at $t=0$ is shown as a dashed black line. The trend we retrieve corresponds to the general expectation that higher $\sigma_{i,0}$ produce higher-energy particles (e.g. Werner & Uzdensky 2021), but with important differences between species and between 2D and 3D runs.
@@
First, we note that both particle species can develop nonthermal populations, with $\sigma_{i,0}=10$ producing evident nonthermal tails extending for up to 2 decades in both 2D and 3D. However, the slope of the nonthermal tail differs depending on dimensionality, with 3D runs producing markedly less steep tails. Reference slopes of $\gamma^{-1.5}$, $\gamma^{-2}$, and $\gamma^{-2.5}$ are shown to guide the eye; this broadly aligns with previous findings for 3D reconnection (Zhang et al. 2021a, 2023). In these runs, ions are initially nonrelativistic and, for the highest $\sigma_{i,0}$, can reach maximum Lorentz factors of $\sim 10$100100; electrons are already relativistic at initialization, and their energy gain is such that $\gamma$ can increase by over 2 orders of magnitude (up to $\sim 10^{5}$) for the case with the highest magnetization. The maximum Lorentz factor is roughly comparable in 2D and 3D for both species, but in 3D it is generally smaller by a factor 2–3 for the highest $\sigma_{i,0}$.
@@ 4 Discussion
We find that low-$\sigma_{i,0}$ runs proceed initially in a 2D-like fashion; tearing modes grow first, creating elongated flux tubes (plasmoids in 2D) which can be destroyed by violent MHD kinking. Short-wavelength DKI modes are relatively unimportant, and overall this results in the dissipation of more magnetic energy than in 2D. At moderate $\sigma_{i,0}=0.4$, DKI modes become strong enough to drive additional magnetic-energy destruction without impeding tearing modes. The latter are still free to develop and break the sheet, creating large-scale MHD-kink-prone flux ropes. Essentially, up until $\sigma_{i,0}=0.4$ we observe no strong competition between tearing and DKI modes, and their combined action releases a much larger fraction of magnetic energy in 3D with respect to 2D (about 45% versus 27%, see Fig. 1aabb). In high-$\sigma_{i,0}$ ($\geq 5$) runs, instead, DKI grows rapidly and actively impedes the thinning of the sheet and the development of tearing. Large-scale flux ropes develop less freely, and instead the current sheet quickly becomes a thick and chaotic structure where magnetic fields are destroyed much more slowly [^5] While Zhang et al. (2021a) reported an observed slowdown of pair-plasma reconnection in 3D vs. 2D, the slowdown observed in our simulations is much more significant.. As a result, a smaller fraction of magnetic energy is dissipated in 3D than in 2D (27% versus 33%). In this situation, particle acceleration is roughly twice less efficient in 3D than in equivalent 2D runs where kinking is absent.
While several points of what reported above deserve more exploration (for example, it remains unclear why $\sigma_{i,0}=0.4$ represents a threshold for the inversion in the energy-dissipation trend), these findings contradict the standard picture in which 3D reconnection dissipates more magnetic energy than in 2D (e.g. Werner & Uzdensky 2021; Zhang et al. 2021b; Werner & Uzdensky 2024). In particular, our results describe a scenario in which reconnecting plasma gathers in slowly dissipating current sheets, which are kept thick and chaotic by DKI. Such a very thick current sheet may even be described as two separate sheets, where upstream, cold- and magnetized-plasma conditions meet an internal hot, lowly magnetized plasma state. In this situation, standard expectations for symmetric Harris-sheet reconnection may not be applicable. Our high-$\sigma_{i,0}$ results are especially relevant for astrophysical scenarios such as compact-object magnetospheres and coronae, where ultramagnetized plasma could be undergoing reconnection, potentially powering high-energy flares. This may have implications for comparing the time scales of reconnection with flaring activity observed from high-energy sources, although quantifying the impact of the current-sheet structure on flare dynamics requires dedicated efforts to be pursued elsewhere.